# Learning-Augmented Robust Algorithmic Recourse

### v2

@@ 4 Experiments
In this section, we describe the datasets, implementation details, and present our findings.
@@
Datasets We experiment on synthetic and real-world data. For the synthetic dataset, we follow Upadhyay et al. [70] to generate 1000 data points in 2-d. For each data point, we first sample a label $y\in\{0,1\}$ uniformly at random. We then sample the instance corresponding to this label from a Gaussian distribution $\mathcal{N}(\mu_{y},\Sigma_{y})$. We set $\mu_{0}=[-2,-2]$, $\mu_{1}=[+2,+2]$, and $\Sigma_{0}=\Sigma_{1}=0.5\mathbb{I}$. See Figure 1(a) in [70]. We also use two real datasets: (a) The German Credit dataset [34], which consists of 1000 data points, each with 7 features containing information about a loan applicant (age, marital status, income, and credit duration), and binary labels good (1) or bad (0) determine the creditworthiness. (b) The Small Business Administration dataset [52], which contains the small business loans approved by the State of California from 1989 to 2004. The dataset includes 1159 data points, each with 28 features containing information about the business (business category, zip code, and number of jobs created by the business), and the binary labels indicate whether the small business has defaulted on the loan (0) or not (1). For real-world data, we normalize the features. We use the datasets to learn the initial model $\theta_{0}$. See Appendix D.7 for experiments on a larger dataset.
@@
Implementation Details We use 5-fold cross-validation in all experiments: 4 folds to train the initial model $\theta_{0}$ and the remaining fold to compute recourse. Recourse is only computed for instances with label (0) under $\theta_{0}$, and we report averages over folds and test instances in all experiments. We used logistic regression as our linear model and trained it using Scikit-Learn. As our non-linear model, we used a 3-level neural network with 50, 100, and 200 nodes in each successive layer (same as [70]). The network uses ReLU activation functions, binary cross-entropy loss, and Adam optimizer, and is trained for 100 epochs using PyTorch. See Appendix D.1 for additional implementation details.
@@
To generate recourse, we implemented Algorithm 1 with different $\beta$ values. For non-linear models, we first approximate them locally with LIME [64] (same as [70]). We used the code from [70] and [56] for ROAR and RBR’s implementation as baselines. We next describe our parameter choices. We use binary cross-entropy as the loss function $\ell$ and $L^{1}$ distance as the cost function $c$. We measure closeness using the $L^{\infty}$ norm for model parameters. In Section 4.1, we follow the same procedure as in [70] for selecting $\alpha$ and $\lambda$. We fix an $\alpha$ and greedily search for $\lambda$ that maximizes the recourse validity under the original model $\theta_{0}$. We study the effect of varying $\alpha$ and $\lambda$ in Section 4.2 and Appendix D.3. In each experiment, we specify how $\hat{\theta}$ is selected. See our code and Appendix D.1 for more details.
@@ 4.1 Findings: Learning-Augmented Setting
Our algorithm Pareto dominates ROAR, simultaneously achieving better robustness and consistency for all $\beta$ values. Furthermore, at $\beta=1$, the robustness of ROAR’s recourses is substantially worse compared to ours (e.g., ROAR achieves a robustness of 17.64 on the neural network models learned on synthetic data (Figure 1(b))). In addition, ROAR is much slower than our algorithm (See Section 4.2). We generally observe that the poor consistency of the robust solutions can be decreased substantially with a small increase in robustness. This is also true for decreasing the robustness of the consistent solutions, albeit to a smaller degree.
@@
In this section, we study how prediction error can affect the recourse quality. In particular, for each dataset-model pair, we first compute the correct prediction $\hat{\theta}_{*}$, corresponding to the realized future model (by either shifting the data or temporal changes in data collection [70]). We then generate incorrect predictions by perturbing each coordinate of the correct prediction by different values in $\{+\epsilon,-\epsilon,+2\epsilon,-2\epsilon\}$, corresponding to the amount of error. The $\epsilon$ values depend on the dataset and model and are chosen to ensure that all the predictions are in $\Theta_{\alpha}$ for $\alpha=1$.
@@
In Figure 2, the rows and columns correspond to different datasets and models. In each sub-figure, curves show the smoothness of Algorithm 1 for different predictions as a function of $\beta$. There are 5 lines in each subfigure corresponding to the correct prediction and the perturbations. Focusing on $\beta=0$, we observe that smoothness increases sharply as a function of the prediction “error”, since the algorithm fully trusts the (incorrect) prediction. As $\beta\to 1$, the smoothness of all predictions converges to the same value, since the algorithm disregards them. In some cases, this convergence occurs at smaller $\beta$ values (Figure 2(e)), but other cases require $\beta$ to be very close to 1 (Figure 2(a)). While the smoothness monotonically increases with $\beta$ when using the correct prediction, using incorrect predictions results in interesting non-monotone behavior and even leads to improved performance compared to using the correct prediction (Figure 2(a)). In Appendix D.5, we provided baselines using a variant of ROAR, though the generated recourses generally have higher smoothness.
@@ 4.2 Findings: Computing Robust Recourse
By setting $\beta=1$, our algorithm can be used to compute robust recourse. In this section, we perform a more detailed comparison with ROAR by breaking down the robustness to understand the effect of each term in Equation (1). The first term is a proxy for validity and the second term is the cost of modifying $x_{0}$. To compute validity, as opposed to computing a possibly different model for each instance, as is done till now, we compute a single model for the entire dataset. Following prior work [70, 56], this model is obtained by training a model on a shifted version of the datasets. We call the validity with respect to this new model as future validity. See Appendix D.1.
@@
Figure 3 indicates that our algorithm often Pareto dominates ROAR, i.e., it achieves higher validity with a smaller cost. While the validity of our algorithm often approaches 1 for smaller $\lambda$ values, this is not the case for ROAR. In addition, the validity is lower for both approaches when using neural network models. Finally, consistent with Pawelczyk et al. [59]’s observation, the cost of recourse can increase significantly for validity values close to 1. In Section D.6, we provide further comparisons with RBR [56].
@@ 5 Conclusion and Discussion
The robust recourse literature has focused on many aspects: different model classes, costs, model changes, and optimization approaches resulting in formulations that require different solutions (see [35]). Furthermore, adopting the learning-augmented framework introduces other modeling aspects such as prediction, definitions of robustness or consistency, and measuring the trade-off between them. We initiated the study of the learning-augmented robust recourse and followed the assumptions in the closest prior work to allow for a direct comparison. We highlight several extensions.
First, while our framework can handle customizable weights for different inputs, using any norm as the cost function implies that the features can be modified independently. Some prior work focuses on considering these dependencies and also actionability of recourse [39, 37, 57]. We leave these as future work. Our notion of robustness and consistency measures the performance of the algorithm against optimal solutions in an additive manner, similar to regret [16]. This comparison can also be made multiplicatively, similar to competitive ratio [53], which we leave as future work. Moreover, studying weaker notions of model change, such as $L^{1}$$L^{2}$ norm [70] or studying alternative formalizations of model change [31] is left for future work. As is common in the learning-augmented framework [1, 53], we assumed the prediction about the updated model is explicitly given. A natural way to compute such a prediction is through performativity [61]. Moreover, in practice, the feedback might be weaker or even noisy [8]. Incorporating these into our framework is an exciting future work direction. Finally, although we did provide non-trivial theoretical results, our focus was on the empirical evaluation of our algorithms and, in line with prior robust recourse work. We leave a potential theoretical analysis to future work.
@@ Acknowledgment
Appendix A Additional Related Work
The emerging literature on the interpretability and explainability of machine learning systems mainly advocates for two main approaches. The first approach aims to build inherently simple or interpretable models such as decision lists [44] or generalized additive models [77, 81]. These approaches provide global explanations for the deployed models. The second approach attempts to explain the decisions of complex black-box models (such as deep neural networks) only on specific inputs [64, 49, 68, 69, 66, 1, 42, 7, 22]. These approaches provide a local explanation of the model and are sometimes referred to as post-hoc explanations.
Recourse is a post-hoc counterfactual explanation that aims to provide the lowest cost modification that changes the prediction for a given input with an undesirable prediction under the current model [76, 71, 48, 63, 67, 59, 26]. Since its introduction, different formulations have been used to model the optimization problem in recourse (see [73] for an overview). Wachter et al. [76] and Pawelczyk et al. [59] considered score-based classifiers and defined modifications to help instances achieve the desired scores. On the other hand, for binary classifiers, Ustun et al. [71] required the modification to result in the desired label. Roughly speaking, the first setting can be viewed as a relaxation of the second setting, and we follow the second formulation in our work.
The follow-up works on the problem study several other aspects such as focusing on specific models such as linear models [71] or decision-trees [38, 12], understanding the setting and its implicit assumptions and implications [6, 72, 23, 25], attainability or actionability [37, 57, 39, 71], imperfect causal knowledge [40], fairness in terms of cost of implementation for different subgroups [28, 30, 33], repeated dynamics [24, 9, 22], temporal data [15] and providing model-agnostic explanations for recourse by computing responsiveness scores [17]. Extending our work to account for these different aspects of recourse is left for future work.
The most closely related work to us is by Upadhyay et al. [70], which initiates the study of robust recourse and proposes an algorithm called RObust Algorithmic Recourse (ROAR). We study the same framework for our robustness analysis. Nguyen et al. [56] proposes a new framework for robust recourse that uses a different objective than ROAR and focuses on data shifts instead of model shifts. They call their framework Robust Bayesian Recourse (RBR) and we refer to their algorithm with the same abbreviation. The RoCourseNet algorithm [29] also provides robust recourse, though a direct comparison with this algorithm is not possible, as it is an end-to-end approach, i.e., it simultaneously optimizes for the learned model and robust recourse while the initial model is fixed in our approach, just like in [70, 56]. Similar to the literature on algorithmic recourse, the literature on robust algorithmic recourse has also studied specialized settings (for different model classes, cost functions, and model changes) and additional desiderata such as fairness [45, 82, 35, 55, 21, 54]. Neither approaches use predictions, though we compare our algorithm with both ROAR and RBR when studying the trade-off between validity and cost of robust recourse in the absence of predictions. See [36] for a recent survey.
In addition to the above works, Pawelczyk et al. [60] studies how model updates due to the “right to be forgotten" can affect recourse validity. Dominguez-Olmedo et al. [20] showed that minimum cost recourse solutions are provably not robust to adversarial perturbations in the model and then presented robust recourse solutions for linear and differentiable models. Black et al. [13] observe that recourse in deep models can be invalid by small perturbations and suggest that the model’s Lipschitzness at the counterfactual point is the key to preserving validity. Hamman et al. [31] proposed a new notion of model change, which they coin “naturally-occurring" model change, and provide recourse with theoretical guarantees on the validity of the recourse.
Recourse is also closely related to adversarial training or robust machine learning [51, 79, 3]. Pawelczyk et al. [58] studied the connections between various recourse formulations and their analogs in the robust machine-learning literature. In fact, ROAR [70] builds on gradient-based methods originally developed for adversarial training. The convergence of such algorithms has been studied extensively under various assumptions (see, e.g., [78]). Theoretical guarantees for adversarial training typically rely on specific data distributions and model classes. For instance, with data drawn from a mixture of Gaussians, the optimal robust classifiers are linear [46]. Moreover, Awasthi et al. [4] propose an algorithm for learning robust linear classifiers under the realizability assumption and establish hardness results for robust learning of degree-2 polynomial threshold functions. Beyond linear models, theoretical analyses have also been extended to other classes such as decision trees [74, 75]. To our knowledge, however, none of these algorithms directly address our setting.
The learning-augmented framework is applied to a wide variety of settings, aiming to provide a refined understanding of the performance guarantees that are achievable beyond the worst case. Unlike approaches in robust optimization that utilize uncertainty sets [10], the learning augmented framework does not impose any assumptions on the quality of the prediction. One of the main application domains is the design of algorithms (e.g., online algorithms [50, 62]), but it has also been used in data structures design [43], mechanisms interacting with strategic agents [2, 80], and privacy-preserving methods for processing sensitive data [41]. This is already a vast and rapidly growing literature; see [47] for a frequently updated and organized list of related papers.
Appendix B ROAR
@@
For completeness, we provide the details of ROAR [70] in this section. ROAR is inspired by the vast literature on adversarial training (see e.g., [51] and utilizes Danskin’s Theorem [18]) to compute the gradients (with respect to $x^{\prime}$). The pseudocode is provided in Algorithm 3. ROAR requires the function $J$ to be differentiable for $x^{\prime}$. Furthermore, ROAR relies on the ability to compute a maximizer of $J$ for $\theta^{\prime}$. When $J$ is differentiable for $\theta^{\prime}$, a local maximum can be computed with projected gradient ascent.
@@
Appendix C Omitted Details from Section 3.1
C.1 Non-convexity of Robust Recourse for Linear Models
@@
We provide a concrete example that makes it easy to verify the non-convexity of the optimization problem in Equation (3) even for linear models. Consider an instance in one dimension where $x_{0}=[1,1]$ (note that the second dimension is the unchangeable intercept), $\theta_{0}=[0,0]$, $\ell$ is squared loss, $\alpha=0.5$, and $\lambda=1$. For any recourse, $x_{r}=[x,1]$ (note that the intercept cannot change), the worst-case $\theta^{\prime}$ is of the form $0.5\text{sign}(x),-0.5]$ since $\alpha$ is 0.5 and $\theta_{0}$ is 0 in both dimensions. The cost of recourse for $x_{r}$ can be written as $1/\left(e^{0.5x\text{sign}(x)-0.5}\right)^{2}+|x-1|$. Plotting this one-dimensional function proves that this function is not convex.
C.2 Proof of Theorem 3.3
@@
Observation C.1 provides an alternative interpretation of the problem: by choosing a recourse $x^{\prime}$, we suffer a cost $\|x^{\prime}-x_{0}\|_{1}$ and the adversary then chooses a $\theta^{\prime}$ aiming to minimize the value of the inner product $x^{\prime}\cdot\theta^{\prime}$. This implies that for a given $x^{\prime}$, a choice of $\theta^{\prime}$ is not optimal for the adversary unless it minimizes this inner product. Also, it implies that among all choices of $x^{\prime}$ with the same cost $\|x^{\prime}-x_{0}\|_{1}$, the optimal one has to maximize the inner product $x^{\prime}\cdot\theta^{\prime}$ with the adversarially chosen $\theta^{\prime}$. We use this fact to prove that a recourse $x^{\prime}$ is not a robust choice by providing an alternative recourse with the same cost and a greater dot product.
@@
Lemma C.2 shows that for any recourse $x$, an adversarial response that minimizes $x\cdot\theta^{\prime}$ is $\theta^{\prime}=\theta_{0}-\alpha\cdot\operatorname{sgn}(x)$. Our next lemma shows how the adversarial response to the initial point $x_{0}$, (i.e., $\theta_{0}-\alpha\cdot\operatorname{sgn}(x_{0})$) determines the direction toward which each dimension of $x_{0}$ should be changed (if at all).
Lemma C.3.
For any optimal recourse $x_{r}\in\arg\min_{x^{\prime}\in\mathcal{X}}\penalty 10000\ \max_{\theta^{\prime}\in\Theta_{\alpha}}J(x^{\prime},\theta^{\prime})$ and every coordinate $i$, it must be that $x_{r}$ raises the value of the $i$-th dimension only if the adversary’s best response to its original value is positive, and it lowers it only if the adversary’s best response to its original value is negative. Using Lemma C.2, we can formally define this as:
@@
Note that in the $k$-th iteration of the while-loop, line 12 of Algorithm 4 chooses $j_{k}$ so that $j_{k}=\arg\max_{j\in\textsc{Active}}|\theta^{\prime}[j]|$, based on the values of $\theta^{\prime}$ at the beginning of that iteration. As a result, if $\theta^{\prime}$ remains the same throughout the execution of the algorithm (which would happen if $\operatorname{sgn}(x_{r})=\operatorname{sgn}(x_{0})$, i.e., if none of the recourse coordinates changes from positive to negative or vice versa), then the lemma is clearly true. On the other hand, if the recourse “flips signs” for some dimension $i$, i.e., $\operatorname{sgn}(x_{r}[i])\neq\operatorname{sgn}(x_{0}[i])$, this could lead to a change of the value of $\theta^{\prime}[i]$. Specifically, as shown in Lemma C.2 and implemented in line 20 of the algorithm, the adversary changes $\theta^{\prime}[i]$ to $\theta_{0}[i]+\alpha\cdot\operatorname{sgn}(x_{0}[i])$. If that transition causes the sign of $\theta^{\prime}[i]$ to change, then dimension $i$ becomes inactive and the algorithm will not consider it again in the future. If the sign of $\theta^{\prime}[i]$ remains the same, then we can show that its absolute value would drop after this change, so even if it is considered in the future, it would still satisfy the claim of this lemma. To verify that its absolute value drops, assume that $x_{0}[i]>0$, suggesting that the algorithm has so far lowered its value to 0, which would only happen if $\theta_{0}[i]<0$ (otherwise, this change would be decreasing the inner product). Since $x_{0}[i]>0$, the new value of $\theta^{\prime}[i]$ is equal to $\theta_{0}[i]+\alpha$, and since this remains negative, like $\theta_{0}[i]$, we conclude that its absolute value decreased. A symmetric argument can be used for the case where $x_{0}[i]<0$. ∎
@@
Note that since $x^{*}\in\arg\min_{x^{\prime}\in\mathcal{X}}\penalty 10000\ \max_{\theta^{\prime}\in\Theta_{\alpha}}J(x^{\prime},\theta^{\prime})$, it must satisfy Lemma C.3. Also, note that the way that Algorithm 4 generates $x_{r}$ also satisfies the conditions of Lemma C.3 (the choice of $\Delta$ in line 13 would never lead to a recourse of higher cost without improving the inner product), so we can conclude that if $x^{*}$ and $x_{0}$ were to change the same coordinate they would both do so in the same direction, i.e.,

### v3

@@ 4 Experiments
In this section, we describe the datasets, implementation details, and present our findings. Our code and data are available at https://github.com/kshitij-kayastha/LearningAugmentedRobustRecourse.
@@
Datasets We experiment on synthetic and real-world data. For the synthetic dataset, we follow Upadhyay et al. (2021) to generate 1000 data points in 2-d. For each data point, we first sample a label $y\in\{0,1\}$ uniformly at random. We then sample the instance corresponding to this label from a Gaussian distribution $\mathcal{N}(\mu_{y},\Sigma_{y})$. We set $\mu_{0}=[-2,-2]$, $\mu_{1}=[+2,+2]$, and $\Sigma_{0}=\Sigma_{1}=0.5\mathbb{I}$. See Figure 1(a) in (Upadhyay et al., 2021). We also use two real datasets: (a) The German Credit dataset (Hofmann,Hans, 1994), which consists of 1000 data points, each with 7 features containing information about a loan applicant (age, marital status, income, and credit duration), and binary labels good (1) or bad (0) determine the creditworthiness. (b) The Small Business Administration dataset (Min Li and Taylor, 2018), which contains the small business loans approved by the State of California from 1989 to 2004. The dataset includes 1159 data points, each with 28 features containing information about the business (business category, zip code, and number of jobs created by the business), and the binary labels indicate whether the small business has defaulted on the loan (0) or not (1). For real-world data, we normalize the features. We use the datasets to learn the initial model $\theta_{0}$. See Appendix C.7 for experiments on a larger dataset.
@@
Implementation Details We use 5-fold cross-validation in all experiments: 4 folds to train the initial model $\theta_{0}$ and the remaining fold to compute recourse. Recourse is only computed for instances with label (0) under $\theta_{0}$, and we report averages over folds and test instances in all experiments. We used logistic regression as our linear model and trained it using Scikit-Learn. As our non-linear model, we used a 3-layer neural network with 50, 100, and 200 nodes in each successive layer (same as (Upadhyay et al., 2021)). The network uses ReLU activation functions, binary cross-entropy loss, and Adam optimizer, and is trained for 100 epochs using PyTorch. See Appendix C.1 for additional implementation details.
@@
To generate recourse, we implemented Algorithm 1 with different $\beta$ values. For non-linear models, we first approximate them locally with LIME (Ribeiro et al., 2016) (same as (Upadhyay et al., 2021)). We used the code from (Upadhyay et al., 2021) and (Nguyen et al., 2022) for ROAR and RBR’s implementation as baselines. We next describe our parameter choices. We use binary cross-entropy as the loss function $\ell$ and $L^{1}$ distance as the cost function $c$. We measure closeness using the $L^{\infty}$ norm for model parameters. In Section 4.1, we follow the same procedure as in (Upadhyay et al., 2021) for selecting $\alpha$ and $\lambda$. We fix an $\alpha$ and greedily search for $\lambda$ that maximizes the recourse validity under the original model $\theta_{0}$. We study the effect of varying $\alpha$ and $\lambda$ in Section 4.2 and Appendix C.3. In each experiment, we specify how $\hat{\theta}$ is selected. See our code and Appendix C.1 for more details.
@@ 4.1 Findings: Learning-Augmented Setting
Our algorithm Pareto dominates ROAR, simultaneously achieving better robustness and consistency for all $\beta$ values. Furthermore, at $\beta=1$, the robustness of ROAR’s recourses is substantially worse compared to ours (e.g., ROAR achieves a robustness of 17.64 on the neural network models learned on synthetic data (Figure 1(b))). ROAR is also much slower than our algorithm (See Section 4.2). We generally observe that the poor consistency of the robust solutions can be decreased substantially with a small increase in robustness. This is also true for decreasing the robustness of the consistent solutions, albeit to a smaller degree.
@@
In this section, we study how prediction error can affect the recourse quality. In particular, for each dataset-model pair, we first compute the correct prediction $\hat{\theta}_{*}$, corresponding to the realized future model (by either shifting the data or temporal changes in data collection (Upadhyay et al., 2021)). We then generate incorrect predictions by perturbing each coordinate of the correct prediction by different values in $\{+\epsilon,-\epsilon,+2\epsilon,-2\epsilon\}$, corresponding to the amount of error. The $\epsilon$ values depend on the dataset and model and are chosen to ensure that all the predictions are in $\Theta_{\alpha}$ for $\alpha=1$.
@@
In Figure 2, the rows and columns correspond to different datasets and models. In each sub-figure, curves show the smoothness of Algorithm 1 for different predictions as a function of $\beta$. There are 5 lines in each subfigure corresponding to the correct prediction and the perturbations. Focusing on $\beta=0$, we observe that smoothness increases sharply as a function of the prediction “error”, since the algorithm fully trusts the (incorrect) prediction. As $\beta\to 1$, the smoothness of all predictions converges to the same value, since the algorithm disregards them. In some cases, this convergence occurs at smaller $\beta$ values (Figure 2(e)), but other cases require $\beta$ to be very close to 1 (Figure 2(a)). While the smoothness monotonically increases with $\beta$ when using the correct prediction, using incorrect predictions results in interesting non-monotone behavior and even leads to improved performance compared to using the correct prediction (Figure 2(a)). In Appendix C.5, we provided baselines using a variant of ROAR, though the generated recourses generally have higher smoothness.
@@ 4.2 Findings: Computing Robust Recourse
By setting $\beta=1$, our algorithm can be used to compute robust recourse. In this section, we perform a more detailed comparison with ROAR by breaking down the robustness to understand the effect of each term in Equation (1). The first term is a proxy for validity and the second term is the cost of modifying $x_{0}$. To compute validity, as opposed to computing a possibly different model for each instance, as is done till now, we compute a single model for the entire dataset. Following prior work (Upadhyay et al., 2021; Nguyen et al., 2022), this model is obtained by training a model on a shifted version of the datasets. We call the validity with respect to this new model as future validity. See Appendix C.1.
@@
Figure 3 indicates that our algorithm often Pareto dominates ROAR, i.e., it achieves higher validity with a smaller cost. While the validity of our algorithm often approaches 1 for smaller $\lambda$ values, this is not the case for ROAR. In addition, the validity is lower for both approaches when using neural network models. Moreover, consistent with Pawelczyk et al. (2023a)’s observation, the cost of recourse can increase significantly for validity values close to 1.
Finally, the LIME approximation can be unreliable, affecting the quality of the trade-off generated by our algorithm. In Appendix C.6, we provide further comparisons with RBR (Nguyen et al., 2022), an approach designed to address this shortcoming. We observe that our algorithm performs similarly to or better than RBR even when the quality of approximation is not perfect, and only when this quality is extremely low (German dataset) does its performance become dominated by RBR. Even in such cases, our algorithm still achieves a high validity (while requiring a higher cost). See Appendix C.6 for more details.
@@ 5 Conclusion and Discussion
The robust recourse literature has focused on many aspects: different model classes, costs, model changes, and optimization approaches resulting in formulations that require different solutions (see (Jiang et al., 2024a)). Furthermore, adopting the learning-augmented framework introduces additional modeling aspects, such as prediction, definitions of robustness and consistency, and measures of their trade-off. We initiated the study of the learning-augmented robust recourse and followed the assumptions in the closest prior work to allow for a direct comparison. We highlight several extensions.
First, while our framework can handle customizable weights for different inputs, using any norm as the cost function implies that the features can be modified independently. Some prior work considers these dependencies and the actionability of recourse (Karimi et al., 2020a; Joshi et al., 2019; Pawelczyk et al., 2020). We leave these as future work. Second, our notion of robustness and consistency measures the performance of the algorithm against optimal solutions in an additive manner, similar to regret (Cesa-Bianchi and Lugosi, 2006). This comparison can also be made multiplicatively, similar to competitive ratio (Mitzenmacher and Vassilvitskii, 2020), which we leave as future work. Third, as is common in the learning-augmented framework (Agarwal et al., 2021; Mitzenmacher and Vassilvitskii, 2020), we assumed the prediction about the updated model is explicitly given. A natural way to compute such a prediction, in practice, is through performativity (Perdomo et al., 2020). For example, the implementation of recourse can cause the distribution shift, and the designer can use this knowledge to form a prediction for the anticipated future model (König et al., 2025). However, this prediction can be imperfect since individuals might not exactly implement the recourse (Fonseca et al., 2023). Moreover, in practice, the feedback might be weaker or even noisy (Bechavod et al., 2022). Incorporating these into our framework is an exciting future work direction. We also leave a potential theoretical analysis or study new robustness frameworks to future work.
@@ Acknowledgment
Appendix A ROAR
@@
For completeness, we provide the details of ROAR [Upadhyay et al., 2021] in this section. ROAR is inspired by the vast literature on adversarial training (see e.g., [Madry et al., 2018] and utilizes Danskin’s Theorem [Danskin, 1967]) to compute the gradients (with respect to $x^{\prime}$). The pseudocode is provided in Algorithm 3. ROAR requires the function $J$ to be differentiable for $x^{\prime}$. Furthermore, ROAR relies on the ability to compute a maximizer of $J$ for $\theta^{\prime}$. When $J$ is differentiable for $\theta^{\prime}$, a local maximum can be computed with projected gradient ascent.
@@
Appendix B Omitted Details from Section 3.1
B.1 Non-convexity of Robust Recourse for Linear Models
Proposition B.1.
Suppose $f$ is a generalized linear model, $c$ is the $L^{1}$ norm, and $\ell(.,1)$ is a convex loss function that is decreasing in its first argument. Then, there exist choices for $\ell$ and $x_{0}$ such that the objective in Equation 3 is non-convex in $x$.
Proof.
@@ B.1 Non-convexity of Robust Recourse for Linear Models
We provide a concrete example that makes it easy to verify the non-convexity of the optimization problem in Equation (3) even for linear models. Consider an instance in one dimension where $x_{0}=[1,1]$ (note that the second dimension is the unchangeable intercept), $\theta_{0}=[0,0]$, $\ell$ is squared loss, $\alpha=0.5$, and $\lambda=1$. For any recourse, $x_{r}=[x,1]$ (note that the intercept cannot change), the worst-case $\theta^{\prime}$ is of the form $0.5{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\operatorname{sgn}}}(x),-0.5]$ since $\alpha$ is 0.5 and $\theta_{0}$ is 0 in both dimensions. The cost of recourse for $x_{r}$ can be written as $1/\left(e^{0.5x{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\operatorname{sgn}}}(x)-0.5}\right)^{2}+|x-1|$. Plotting this one-dimensional function proves that this function is not convex. ∎
B.2 Proof of Theorem 3.3
@@
Observation B.2 provides an alternative interpretation of the problem: by choosing a recourse $x^{\prime}$, we suffer a cost $\|x^{\prime}-x_{0}\|_{1}$ and the adversary then chooses a $\theta^{\prime}$ aiming to minimize the value of the inner product $x^{\prime}\cdot\theta^{\prime}$. This implies that for a given $x^{\prime}$, a choice of $\theta^{\prime}$ is not optimal for the adversary unless it minimizes this inner product. Also, it implies that among all choices of $x^{\prime}$ with the same cost $\|x^{\prime}-x_{0}\|_{1}$, the optimal one has to maximize the inner product $x^{\prime}\cdot\theta^{\prime}$ with the adversarially chosen $\theta^{\prime}$. We use this fact to prove that a recourse $x^{\prime}$ is not a robust choice by providing an alternative recourse with the same cost and a greater dot product.
@@
Lemma B.3 shows that for any recourse $x$, an adversarial response that minimizes $x\cdot\theta^{\prime}$ is $\theta^{\prime}=\theta_{0}-\alpha\cdot\operatorname{sgn}(x)$. Our next lemma shows how the adversarial response to the initial point $x_{0}$, (i.e., $\theta_{0}-\alpha\cdot\operatorname{sgn}(x_{0})$) determines the direction toward which each dimension of $x_{0}$ should be changed (if at all).
Lemma B.4.
For any optimal recourse $x_{r}\in\arg\min_{x^{\prime}\in\mathcal{X}}\penalty 10000\ \max_{\theta^{\prime}\in\Theta_{\alpha}}J(x^{\prime},\theta^{\prime})$ and every coordinate $i$, it must be that $x_{r}$ raises the value of the $i$-th dimension only if the adversary’s best response to its original value is positive, and it lowers it only if the adversary’s best response to its original value is negative. Using Lemma B.3, we can formally define this as:
@@
Note that in the $k$-th iteration of the while-loop, line 12 of Algorithm 4 chooses $j_{k}$ so that $j_{k}=\arg\max_{j\in\textsc{Active}}|\theta^{\prime}[j]|$, based on the values of $\theta^{\prime}$ at the beginning of that iteration. As a result, if $\theta^{\prime}$ remains the same throughout the execution of the algorithm (which would happen if $\operatorname{sgn}(x_{r})=\operatorname{sgn}(x_{0})$, i.e., if none of the recourse coordinates changes from positive to negative or vice versa), then the lemma is clearly true. On the other hand, if the recourse “flips signs” for some dimension $i$, i.e., $\operatorname{sgn}(x_{r}[i])\neq\operatorname{sgn}(x_{0}[i])$, this could lead to a change of the value of $\theta^{\prime}[i]$. Specifically, as shown in Lemma B.3 and implemented in line 20 of the algorithm, the adversary changes $\theta^{\prime}[i]$ to $\theta_{0}[i]+\alpha\cdot\operatorname{sgn}(x_{0}[i])$. If that transition causes the sign of $\theta^{\prime}[i]$ to change, then dimension $i$ becomes inactive and the algorithm will not consider it again in the future. If the sign of $\theta^{\prime}[i]$ remains the same, then we can show that its absolute value would drop after this change, so even if it is considered in the future, it would still satisfy the claim of this lemma. To verify that its absolute value drops, assume that $x_{0}[i]>0$, suggesting that the algorithm has so far lowered its value to 0, which would only happen if $\theta_{0}[i]<0$ (otherwise, this change would be decreasing the inner product). Since $x_{0}[i]>0$, the new value of $\theta^{\prime}[i]$ is equal to $\theta_{0}[i]+\alpha$, and since this remains negative, like $\theta_{0}[i]$, we conclude that its absolute value decreased. A symmetric argument can be used for the case where $x_{0}[i]<0$. ∎
@@
Note that since $x^{*}\in\arg\min_{x^{\prime}\in\mathcal{X}}\penalty 10000\ \max_{\theta^{\prime}\in\Theta_{\alpha}}J(x^{\prime},\theta^{\prime})$, it must satisfy Lemma B.4. Also, note that the way that Algorithm 4 generates $x_{r}$ also satisfies the conditions of Lemma B.4 (the choice of $\Delta$ in line 13 would never lead to a recourse of higher cost without improving the inner product), so we can conclude that if $x^{*}$ and $x_{0}$ were to change the same coordinate they would both do so in the same direction, i.e.,